OpenStudy (anonymous):

what is the distance from _7 to 15 on a number line

OpenStudy (princeharryyy):

-7 to 0 =>7 0 to 12 => 12 , total 7 + 12 =19

OpenStudy (princeharryyy):

@alex122is

OpenStudy (princeharryyy):

sorry 0 to 15 => 15 so answer = 7 +15 = 22 @alex122is

OpenStudy (anonymous):

its that easy lol

OpenStudy (princeharryyy):

yup it is @alex122is

OpenStudy (princeharryyy):

nything else

OpenStudy (anonymous):

is the square root of 11 an irrational number ??

OpenStudy (anonymous):

can you factore the two cubed

OpenStudy (anonymous):

no explanation im in a hurry

OpenStudy (princeharryyy):

see every square root is an irrational function if it isn't a perfect square or if it can't be written in the form of p/q .... yes it is irrational

OpenStudy (anonymous):

c^3-T^3

OpenStudy (princeharryyy):

yup

OpenStudy (anonymous):

yea thank you hurry plsss

OpenStudy (princeharryyy):

(c-t)(c^2 + t^2 +ct)

OpenStudy (princeharryyy):

next !!! @alex122is

OpenStudy (anonymous):

81c^2-4d^2

OpenStudy (anonymous):

factor

OpenStudy (anonymous):

u have a minute

OpenStudy (princeharryyy):

(9c-2d)(9c+2d)

OpenStudy (princeharryyy):

next @alex122is

OpenStudy (anonymous):

omg i passed bc of you

OpenStudy (anonymous):

omg thank youuuuuuuuuuuu

OpenStudy (princeharryyy):

u r welcome @alex122is it was fun !!!

OpenStudy (anonymous):

you made my day

OpenStudy (anonymous):

you see im taking an online class for college. and we were supposed to do hw. We had 100problems in the passed two week. the problem was that my account had techinicle difficulties, so i just got my account today. I did 50 propblems and this test.

OpenStudy (princeharryyy):

ohhh !!! dats gud @alex122is

OpenStudy (anonymous):

well now i have a 50 percent in the class :(

OpenStudy (princeharryyy):



OpenStudy (princeharryyy):

u'll soon b at top @alex122is Have a Nice Day !!!! Gud Luck !!!

OpenStudy (anonymous):

thank you. you too <3

OpenStudy (princeharryyy):

if u need help on anyother problem u can always message me ... bieee )))

OpenStudy (princeharryyy):

tnks )))

OpenStudy (anonymous):

ok yea ill add you <3. thank you

OpenStudy (princeharryyy):

))) biee